# Adaptive Self-improvement LLM Agentic System for ML Library Development

### v1

@@ Abstract
ML libraries, often written in architecture-specific programming languages (ASPLs) that target domain-specific architectures, are key to efficient ML systems. However, writing these high-performance ML libraries is challenging because it requires expert knowledge of ML algorithms and the ASPL. Large language models (LLMs), on the other hand, have shown general coding capabilities. However, challenges remain when using LLMs for generating ML libraries using ASPLs because 1) this task is complicated even for experienced human programmers and 2) there are limited code examples because of the esoteric and evolving nature of ASPLs. Therefore, LLMs need complex reasoning with limited data in order to complete this task. To address these challenges, we introduce an adaptive self-improvement agentic system. In order to evaluate the effectiveness of our system, we construct a benchmark of a typical ML library and generate ASPL code with both open and closed-source LLMs on this benchmark. Our results show improvements of up to 3.9×3.9$\times$3.9 × over a baseline single LLM.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: We propose an adaptive self-improvement LLM agentic system. Similar to human experiential learning (Kolb, 2014), LLM agents start from their base knowledge and accumulate experiences through parallel sampling. Our adaptive self-improvement learning algorithm filters high-quality answers, stratifies the earned experiences by difficulty, and adaptively selects demonstrations to enhance LLM agents.
@@
Each DSA needs a corresponding ML library, a collection of ML operators written in its ASPL, before programmers can effectively use the DSA to accelerate ML applications. ML library development is challenging because it requires expertise in both ML algorithms and the target ASPL. Essentially, library development is a generation process that composes low-level ASPL primitives into high-level ML operators (Dong et al., 2024; Ye et al., 2025).
ML library development using ASPLs requires complex reasoning while minimizing data requirements. ML libraries are developed simultaneously as the chip is manufactured to meet the production timeline (Villa et al., 2021). This library–chip co-design process creates such a tight timeline that ASPLs have only a limited number of code examples. Moreover, this task is complicated even for experienced human programmers. For example, the publication of FlashAttention-3 (Shah et al., 2024) lagged behind the release of the H100 by two years. Directly adapting FlashAttention-2 (Dao, 2023) from A100 to H100 GPU witnessed a 47% performance drop (Spector et al., 2024).
@@
To address these limitations, we design an adaptive self-improvement learning algorithm integrated with an agentic system organization. This approach not only produces a self-improving agentic system to assist humans but also generates high-quality ML operators that can be leveraged by other systems. We show our system in Figure 1. Our techniques create a self-improvement cycle: LLM agents evolve through earned experiences and these evolved agents can earn more experiences. This self-improvement cycle is fully automated, involving no human experts beyond the ASPL designers themselves, who initially tell the models how to use the ASPL primitives.
Inspired by curriculum learning (Bengio et al., 2009), our algorithm prioritizes hard-earned experiences gained from challenging task completion. When these hard-earned experiences are used up, the algorithm adaptively increases the number of demonstrations by mixing experiences earned from less challenging tasks. Section 6.1 shows that hard-earned experiences improve LLM agents more efficiently than mixed ones. Mixed experiences, while sometimes diluting demonstrations, can help agents overcome learning obstacles and complete more tasks. As a byproduct, the algorithm adaptively increases test-time compute on challenging tasks until they are finished or the data is used up.
To emulate the library–chip co-design process, we choose Streaming Tensor Programs (STeP) as the target ASPL for library generation. STeP is an emerging ASPL designed for next-generation reconfigurable dataflow architectures (Prabhakar et al., 2017), a family of DSAs for AI (Prabhakar & Jairath, 2021; Chen et al., 2023; Prabhakar et al., 2024). The only public document of STeP is a non-archival three-page workshop publication (Sohn et al., 2024a), which defines its semantics without any code examples or execution environments. Therefore, STeP programs do not exist in the training corpus of any LLM.
@@
Putting all these together, our system solves up to 96% of the tasks in our benchmark and achieves up to 3.9×3.9$\times$3.9 × improvements over a baseline single LLM, as shown in Figure 2. The contributions of this paper are: (1) an adaptive self-improvement learning algorithm that enables LLM agents to continuously construct ML libraries through adaptive experience-driven evolution; (2) an end-to-end agentic system that uses adaptive self-improvement to develop an ML library for STeP, an ASPL for a next-generation AI accelerator; (3) a complete evaluation of the adaptive self-improvement learning algorithm and the integrated agentic system on a realistic benchmark constructed from common ML operators.
Figure 3: We benchmarked 8 groups of ML operators for a common LLM layer (Jiang et al., 2024). They can be categorized into dynamic, static matrix, and static vector operators where parentheses indicate the number of tasks associated with each group of ML operators.
@@ 2 Background
In this section, we provide background on how DSAs are programmed through their ASPLs, describe how these ASPLs are used to create end-user ML libraries, and identify the key challenges of this library generation process. We also establish STeP as the target ASPL to explore LLM techniques for ML library development. Key concepts related to the background are listed in Table 5 for reference.
@@ 2.2 ML library development using ASPLs
ML libraries developed in ASPLs face portability challenges because ASPLs rapidly evolve to align with DSA updates to meet the demand of growing ML workloads. For example, the matrix multiplication units on NVIDIA GPUs and their corresponding MMA instructions have been updated every generation since their first introduction (NVIDIA, 2017). Consequently, every library function that uses MMA instructions must be rewritten in a new ASPL per generation. Moreover, ML libraries need to be shipped at the same time as the chip. In this case, library development costs are no longer negligible.
@@
Figure 4: STeP is an ASPL for next-generation RDAs. (a) diagrams a STeP program for a simplified MoE module. More details can be found in Section A.1. (b) illustrates the streaming dataflow execution of (a) on an RDA where “MU” stands for memory unit, and “CU” stands for compute unit. (c) is the SN40L, a deployed RDA chip.
@@ 2.3 STeP for next-generation RDA
We chose STeP as our target ASPL due to its potential for better efficiency and its status as a research prototype ASPL. STeP’s efficiency potential stems from its role as the ASPL for next-generation RDAs, which have emerged as a promising alternative to GPUs. The SN40L, a deployed RDA implementation shown in Figure 4(c), demonstrates record-breaking inference speeds for the Llama 3.1 405B model (SambaNova, 2024). Although STeP does not yet have a path to a fabricated chip, developing ML libraries in STeP still presents similar challenges as other ASPLs. Writing STeP programs requires complex reasoning about streaming dataflow execution, and our work began without any existing executable STeP programs to reference.
Like other ASPLs, STeP’s operational semantics are expressed by primitives and specialized functions. Specifically, STeP primitives describe different stream token manipulation strategies, and STeP specialized functions express different configurations of RDA units. STeP adds streaming to the conventional dataflow execution model of RDAs as shown in Figure 4(b). Streaming dataflow unifies both data values and control signals as stream tokens, baking the control flow directly into the data, for better dynamism. The streaming dataflow execution model in STeP is inspired by parallel patterns and array programming (Hsu et al., 2023; Rucker et al., 2024). A stream can be consumed by at most one primitive because of the queueing nature of dataflow (Zhang et al., 2021), which is called an affine type constraint in programming language theory (Wikipedia, 2024). The affine type constraint is a global property of the program since it counts the usage of a variable in the whole program.
@@
0:  $\mathcal{X}$: task set, $m$: adaptive granularity
0:  $\theta$: LLM agentic system, $r$: reward from verifier, $\sigma$: filter function, $\beta$: selection function
1:  $\mathcal{D}\leftarrow\emptyset$
2:  $t\leftarrow 0$triangleright iteration
@@
4:     $\mathcal{E}\leftarrow\beta(\mathcal{D},m)$triangleright stratification
@@
6:        $d_{j}\leftarrow[e_{0},e_{1},...e_{j}]$triangleright selection
@@
12:           $\mathcal{D}\leftarrow\mathcal{D}\cup\sigma(\mathcal{B}_{t},\mathcal{C}_{t},\mathcal{S}_{t})$triangleright filtering
@@ 3 Adaptive self-improvement learning
Our adaptive self-improvement learning evolves LLM agentic systems with data generated by themselves. This algorithm parallel samples the agentic system for correct answers with the success rate, filters high-quality correct answers, stratifies the earned experiences, and adaptively updates demonstrations until all the tasks are solved or the demonstrations are used up. Algorithm 1 shows the complete algorithm for adaptive self-improvement learning. As a byproduct, the algorithm adaptively assigns more test-time compute to harder tasks because it excludes a task from the task set after completion. Figure 5 shows a running example, where only the unfinished tasks are fed to the agentic system. This algorithm is independent of the specific organization of the agentic system, as shown in Figure 1.
@@ 3.1 Filtering high-quality answers
The filter function, $\sigma$, collects earned experience $\mathcal{D}$ by filtering one answer for each newly solved task in $\mathcal{S}_{t}$ and records the success rate of this answer. $\sigma$ first groups the correct answers $\mathcal{B}_{t}$ by the isomorphic abstract syntax tree (Knuth, 1968). Then, it randomly selects one answer from each isomorphic group. After that, it selects the one with the minimal length of pure code (excluding comments and empty lines) from these canonical representatives as the final answer for the task in $\mathcal{S}_{t}$. It also stores the success rate from $\mathcal{C}_{t}$ for $\beta$. The minimal length selection follows the Minimum Description Length principle for higher information density (Rissanen, 1978). On the other hand, shorter text might lose chain-of-thought comments (Wei et al., 2022b). To balance these two contradicting intuitions, we introduce randomness to the selection of representative answers for each isomorphic group.
@@ 3.2 Stratification and selection
The selection function $\beta$ stratifies the earned experiences $\mathcal{D}$ by binning them into $m$ levels of difficulty and demonstrations $d_{j}$ are selected incrementally from the stratified experiences $\mathcal{E}$. We define the difficulty as the opposite of the success rate following (Lightman et al., 2023). $\beta$ sorts $\mathcal{D}$ in ascending order of success rate and then bins the tasks as evenly as possible to get the boundaries of each bin. Then tasks are rebinned using these boundary values. This selection strategy can cause repetitive steps as exemplified by the dash line circles (Iter 4&5 in Figure 7(a)) when the newly finished tasks are easier than the demonstrations. Our algorithm keeps these repetitive steps instead of avoiding them because the tasks with low success rates have a better chance of getting one correct answer with the number of samples doubled. For example, Iter 4 performs better than Iter 3 in Figure 7(b) with the same demonstrations. This method can also cause later iterations to have fewer tokens but with higher quality than the previous iteration (Iter 2&3 in Figure 7(a)) when the boundary value crosses two bins and the newly finished tasks are easier.
@@ 3.3 Discussion
To use this technique in a continuous learning setting, users can add new tasks to the initial $\mathcal{X}$ and apply Algorithm 1. This approach, however, does not guarantee success and might exceed context window constraints. If the task set $\mathcal{X}$ is finite, as in our case, the algorithm will terminate.
@@ 4 Agentic system organization
In this section, we introduce a specific agentic system organization tailored for ML library development using STeP as shown in Figure 6. It includes LLM agents, a code generator, and verifiers with a structural intermediate representation as the interface between users and system components.
@@ 4.1 LLM agents
As shown in Figure 6(a), each single LLM is designed for a purpose assigned by the domain expert through the base prompt. The base prompt contains the task description, base knowledge, and demonstrations. The prompt composer chains the base prompt and task-specific input, which is then fed into a configured LLM serving API.
We design two agents, a proposer and a guardian. As explained in Section 2.3, the affine type constraint in STeP is a global property that requires thinking back and forth beyond step-by-step reasoning, which is challenging for the causal generation of LLMs. Therefore, we design a guardian agent to check and correct the affine type error globally. We exclude the demonstration tasks in the base prompt of the agents from the benchmark to avoid the model directly copying the answer.
@@ 4.2 Code generator
After the LLM agents, a code generator takes in the generated implementation and outputs a self-contained pytestable Python script. With this code generator, LLM agents only need to output implementations without other helper code. We embed the STeP specifications written in natural language to Python (as exemplified in Figure 21 to Figure 22). Beneath this Python frontend is a functional simulator that calculates the result of STeP programs. Since the essence of each ASPL is the programming abstraction (semantics) instead of its syntax (Liskov, 2011), we choose to prompt LLMs with their familiar Python syntax. In this way, LLMs can focus more on reasoning about the STeP programming abstraction without being distracted by alien syntax.
@@ 4.3 Verifier
Fast verification is vital because it bottlenecks adaptive self-improvement learning. ASPLs further increase this complexity by requiring a simulator for the library–chip co-design process. Simulating STeP as a general dataflow system in Python would be slow because of its high dynamism (Zhang et al., 2024a). Since we only focus on functional correctness in this work, we meticulously limit the level of dynamism to the degree that ML operators require. Consequently, our simulator emulates stream execution using tensor computation with necessary control flows.
Our system organization contains two verifiers. As a result, the reward in Algorithm 1 is $r(x,y)=1$ for task $y$ if answer $x$ passes these two verifiers and $r(x,y)=0$ otherwise. One verifier checks for functional correctness. Users program PyTorch to express their ML operators, which elicits LLMs’ base knowledge of ML algorithms. The verifier compares the execution results of our simulator with the result tensors of the corresponding PyTorch program on a single set of shapes with random input values. The fidelity of our unit test method builds on practice (Jia et al., 2019) and theory (Gulwani & Necula, 2003). The other verifier checks the affine type constraint by performing static analysis on the abstraction syntax tree of the STeP program with the Python ast module (Ronacher, 2008).

### v2

@@ Abstract
ML libraries, often written in architecture-specific programming languages (ASPLs) that target domain-specific architectures, are key to efficient ML systems. However, writing these high-performance ML libraries is challenging because it requires expert knowledge of both ML algorithms and the ASPL. Large language models (LLMs), on the other hand, have shown general coding capabilities. However, challenges remain when using LLMs for generating ML libraries using ASPLs because 1) this task is complicated even for human experts and 2) there are limited code examples due to the esoteric and evolving nature of ASPLs. We present an adaptive self-improvement agentic system that enables LLMs to perform such complex reasoning under limited data by iteratively improving their capability through self-generated experience. In order to evaluate the effectiveness of our system, we construct a benchmark of a typical ML library and generate ASPL code with both open and closed-source LLMs on this benchmark. Our results show improvements of up to $3.9\times$ over a baseline single LLM  [^1] The example code is public at https://github.com/zhang677/PCL-lite.
LLM agents, Self-improvement learning, Machine learning library
@@ 1 Introduction
Figure 1: We propose an adaptive self-improvement LLM agentic system. LLM agents start from their base knowledge and accumulate experiences through parallel sampling. Our adaptive self-improvement learning algorithm filters high-quality answers, stratifies the earned experiences by difficulty, and adaptively selects demonstrations to enhance LLM agents.
@@
Each DSA needs a corresponding ML library, a collection of ML operators written in its ASPL, before programmers can effectively use the DSA to accelerate ML applications. ML library development is challenging as it requires expertise in both ML algorithms and the target ASPL. Essentially, library development is a generation process that composes low-level ASPL primitives into high-level ML operators (Dong et al., 2024; Ye et al., 2025).
Furthermore, ML library development using ASPLs requires complex reasoning while minimizing data requirements. ML libraries development often occurs in parallel with hardware manufacturing to meet production deadlines (Villa et al., 2021). This time-constrained library–chip co-design process leaves limited code examples. Moreover, this task is complicated even for expert human programmers. For example, the publication of FlashAttention-3 (Shah et al., 2024) lagged behind the release of the H100 by two years. The challenge is further intensified by the need for ML libraries to co-evolve with new hardware to sustain performance. For instance, directly adapting FlashAttention-2 (Dao, 2023) from the A100 to H100 GPU witnessed a 47% performance drop (Spector et al., 2024).
@@
To address these limitations, we design an adaptive self-improvement learning algorithm integrated with an agentic system organization. This approach not only produces a self-improving agentic system to assist humans but also generates high-quality ML operators that can be leveraged by other systems. We show our system in Figure 1. Similar to human experiential learning (Kolb, 2014), our techniques create a self-improvement cycle: LLM agents evolve through earned experiences and these evolved agents can earn more experiences. This self-improvement cycle is fully automated, involving no human experts beyond the ASPL designers themselves, who initially tell the models how to use the ASPL primitives.
Our algorithm prioritizes hard-earned experiences gained from completing challenging tasks, inspired by curriculum learning (Bengio et al., 2009). When these hard-earned experiences are exhausted, the algorithm adaptively increases the number of demonstrations by incoporating experiences from less challenging tasks. Section 6.1 shows that hard-earned experiences improve LLM agents more efficiently than mixed ones. While mixed experiences may dilute demonstrations, they can help agents overcome learning obstacles and complete more tasks. As a byproduct, the algorithm adaptively increases test-time compute on challenging tasks until they are complete or the data is exhausted.
To emulate the library–chip co-design process, we choose Streaming Tensor Programs (STeP) as our target ASPL for library generation. STeP is an emerging ASPL designed for next-generation reconfigurable dataflow architectures (Prabhakar et al., 2017), a family of DSAs for AI (Prabhakar & Jairath, 2021; Chen et al., 2023; Prabhakar et al., 2024). The only public document of STeP is a non-archival three-page workshop publication (Sohn et al., 2024a), which defines its semantics without any code examples or execution environments. Therefore, STeP programs do not exist in the training corpus of any LLM.
@@
Putting all these together, our system solves up to 96% of the tasks in our benchmark and achieves up to a $3.9\times$ improvement over a baseline single LLM, as shown in Figure 2. The contributions of this paper are: (1) an adaptive self-improvement learning algorithm that enables LLM agents to continuously construct ML libraries through adaptive experience-driven evolution; (2) an end-to-end agentic system that uses adaptive self-improvement to develop an ML library for STeP, an ASPL for a next-generation AI accelerator; (3) a complete evaluation of the adaptive self-improvement learning algorithm and the integrated agentic system on a realistic benchmark constructed from common ML operators.
Figure 3: We benchmarked eight groups of ML operators within a common LLM layer (Jiang et al., 2024). These operators are categorized as dynamic, static matrix, and static vector, with parentheses indicating the number of tasks associated with each group.
@@ 2 Background
In this section, we provide background on how DSAs are programmed through their ASPLs, describe how these ASPLs are used to create end-user ML libraries, and identify key challenges of this library generation process. We also establish STeP as the target ASPL to explore LLM techniques for ML library development. Key concepts related to the background are listed in Table 8 for reference.
@@ 2.2 ML library development using ASPLs
ML libraries developed in ASPLs face portability challenges because ASPLs rapidly evolve to align with DSA updates in new generations to meet the demands of growing ML workloads. For example, the matrix multiplication units on NVIDIA GPUs and their corresponding MMA instructions have been updated every generation of GPU since their introduction (NVIDIA, 2017). Consequently, every library function that uses MMA instructions must be rewritten in a new ASPL (e.g. CUDA with new instructions) per generation. Moreover, ML libraries need to be shipped at the same time as the chip. In this case, library development costs are no longer negligible.
@@
Figure 4: STeP is an ASPL for next-generation RDAs that models streaming dataflow execution. (a) STeP example program for a simplified MoE module. More details can be found in Section A.1. (b) An illustration of the streaming dataflow execution of (a) on an RDA of memory units (MU), and compute units (CU). (c) The SN40L, a deployed RDA chip.
@@ 2.3 STeP for next-generation RDA
We chose STeP as our target ASPL due to its potential for better efficiency and status as a research prototype ASPL. STeP’s efficiency potential stems from its role as an ASPL for next-generation RDAs, which are a promising alternative to GPUs. The SN40L, a deployed RDA implementation shown in Figure 4(c), demonstrates record-breaking inference speeds for the Llama 3.1 405B model (SambaNova, 2024). Although STeP does not yet have a path to a fabricated chip, developing ML libraries in STeP still presents similar challenges as other ASPLs. Writing STeP programs requires complex reasoning about streaming dataflow execution, and our work began without any existing executable STeP programs to reference.
Similar to other ASPLs, primitives and specialized functions compose to form operational semantics in STeP. Specifically, STeP primitives describe different stream token manipulation strategies, and STeP specialized functions express different configurations of RDA units. Inspired by parallel patterns and array programming (Hsu et al., 2023; Rucker et al., 2024), the streaming dataflow execution model in STeP treats data as streams of tokens flowing between computational units in space. STeP extends the conventional dataflow execution model of RDAs by introducing streaming semantics, as shown in Figure 4(b). This approach unifies data values and control signals into stream tokens, embedding control flow directly into the data to enable greater dynamism. A stream can be consumed by at most one primitive because of the queueing nature of dataflow (Zhang et al., 2021), which is called an affine type constraint in programming language theory (Wikipedia, 2024). The affine type constraint is a global property of the program since it counts the usage of a variable in the whole program.
@@
[…]
@@
4:  $\mathcal{E}\leftarrow\beta(\mathcal{D},m)$ $\triangleright$ stratification
@@
6:   $d_{j}\leftarrow[e_{0},e_{1},...e_{j}]$ $\triangleright$ selection
@@
12:    $\mathcal{D}\leftarrow\mathcal{D}\cup\sigma(\mathcal{B}_{t},\mathcal{C}_{t},\mathcal{S}_{t})$ $\triangleright$ filtering
@@ 3 Adaptive self-improvement learning
Our adaptive self-improvement learning evolves LLM agentic systems with data generated by the systems themselves. This algorithm samples the agentic system in parallel for correct answers with success rates, filters high-quality correct answers, stratifies the earned experiences, and adaptively updates demonstrations until all the tasks are solved or the demonstrations are exhausted. The complete algorithm for adaptive self-improvement learning is shown in Algorithm 1. As a byproduct, the algorithm adaptively assigns more test-time compute to harder tasks by excluding tasks from the task set after completion. Figure 5 illustrates a running example in which only the unfinished tasks are fed to the agentic system. Additionally, this algorithm is independent of the specific organization of the agentic system, as shown in Figure 1.
@@ 3.1 Filtering high-quality answers
The filter function, $\sigma$, collects earned experience $\mathcal{D}$ by filtering one answer for each newly solved task in $\mathcal{S}_{t}$ and records the success rate of this answer in $\mathcal{C}_{t}$. $\sigma$ first groups the correct answers $\mathcal{B}_{t}$ by the isomorphic abstract syntax tree (Knuth, 1968). Then, $\sigma$ groups equivalent answers using AST isomorphism and randomly selects one representative from each group. Among these representatives, $\sigma$ chooses the one with the minimal length of pure code (excluding comments and empty lines) as the final answer for the task in $\mathcal{S}_{t}$. $\sigma$ also stores the success rate from $\mathcal{C}_{t}$ for $\beta$. The minimal length selection follows the Minimum Description Length principle for higher information density (Rissanen, 1978). On the other hand, shorter text might lose chain-of-thought comments (Wei et al., 2022b). We purposefully introduce randomness to the selection of representative answers for each isomorphic group to balance these two contradicting intuitions.
@@ 3.2 Stratification and selection
The selection function $\beta$ stratifies the earned experiences $\mathcal{D}$ by binning them into $m$ levels of difficulty and demonstrations $d_{j}$ are selected incrementally from the stratified experiences $\mathcal{E}$. We define the difficulty as the opposite of the success rate following (Lightman et al., 2023). $\beta$ sorts $\mathcal{D}$ in ascending order of success rate and then bins the tasks as evenly as possible to get the boundaries of each bin. Then tasks are rebinned using these boundary values. This selection strategy can cause repetitive steps as exemplified by the dashed-line circles (Iter 4&5 in Figure 7(a)) when the newly finished tasks are easier than the demonstrations. Our algorithm keeps these repetitive steps instead of avoiding them because the tasks with low success rates have a better chance of getting one correct answer with the number of samples doubled. For example, Iter 4 performs better than Iter 3 in Figure 7(b) with the same demonstrations. This method can also cause later iterations to have fewer tokens but with higher quality than previous iterations (Iter 2&3 in Figure 7(a)) when the boundary value crosses two bins and the newly finished tasks are easier.
@@ 3.3 Discussion
Algorithm 1 can also extend to new tasks that typically involve new ML operators and new hardware specialized functions. These can be incorporated into the initial task pool $\mathcal{X}$ and handled using the same adaptive self-improvement learning process by selectively sampling only the new tasks. Each example takes about 500 tokens, so cutting-edge LLMs can handle hundreds more tasks. If the number exceeds the context length, better stratification and selection functions are needed to preserve experience quality within the context window limit. If the task set $\mathcal{X}$ is finite, as in our case, the algorithm will terminate.
@@ 4 Agentic system organization
In this section, we introduce a specific agentic system organization tailored for ML library development using STeP as shown in Figure 6. The agentic system comprises LLM agents, a code generator, and verifiers, with a structural intermediate representation as the interface between users and system components.
@@ 4.1 LLM agents
As shown in Figure 6(a), each single LLM is designed for a specific purpose assigned by the domain expert through the base prompt. The base prompt contains the task description, base knowledge, and demonstrations. The prompt composer chains the base prompt and task-specific input, which is then fed into a configured LLM serving API.
Specifically, we design two agents, a proposer and a guardian. As explained in Section 2.3, the affine type constraint in STeP is a global property that requires thinking back and forth beyond step-by-step reasoning, which is challenging for the causal generation of LLMs. Therefore, we design a guardian agent to check and correct the affine type error globally. We exclude the demonstration tasks in the base prompt of the agents from the benchmark to avoid the model directly copying the answer.
@@ 4.2 Code generator
After the LLM agents, a code generator takes in the generated implementation and outputs a self-contained pytestable Python script. With this code generator, the LLM agents only need to output implementations without other helper code. We embed the STeP specifications written in natural language to Python (see Figure 21 and Figure 22). Beneath this Python frontend is a functional simulator that calculates the result of STeP programs. Since the essence of each ASPL is the programming abstraction (semantics) instead of its syntax (Liskov, 2011), we choose to prompt LLMs with their familiar Python syntax. In this way, LLMs can focus more on reasoning about the STeP programming abstraction without being distracted by less familiar syntax.
@@ 4.3 Verifier
Fast verification is vital because it bottlenecks adaptive self-improvement learning. ASPLs further increase this complexity by requiring a simulator for the library–chip co-design process. Simulating STeP as a general dataflow system in Python would be slow because of its high dynamism (Zhang et al., 2024b). Since we only focus on functional correctness in this work, we meticulously limit the level of dynamism to the degree that ML operators require. Consequently, our simulator emulates stream execution using tensor computation with the necessary control flow.
Our system organization contains two verifiers, and the reward function for our system in Algorithm 1 is $r(x,y)=1$ for task $y$ if answer $x$ passes these two verifiers and $r(x,y)=0$ otherwise. One verifier checks for functional correctness. Users program PyTorch to express their ML operators, which elicits LLMs’ base knowledge of ML algorithms. The verifier compares the execution results of our simulator with the result tensors of the corresponding PyTorch program on a single set of shapes with random input values. The fidelity of our unit test method builds on practice (Jia et al., 2019) and theory (Gulwani & Necula, 2003). The other verifier checks the affine type constraint by performing static analysis on the abstraction syntax tree of the STeP program with the Python ast module (Ronacher, 2008).